OpenStudy (anonymous):

rob can tar a roof in nine hours. one day his friend kayla helped him and it only took 4.74 hrs. how long will it take kayla to do it alone?

OpenStudy (wolf1728):

When 2 people work the formula is: Person1 * Person2 / (Person1 + Person2) rob becomes Person1 and Kayla becomes Person 2 9 * K / (9 + K) = 4.74 Hours 9*K = 4.74*9 + K*4.74 9*K = 42.66 + K*4.74 4.26 *K = 42.66 K = 10.014084507 hours (Kayla) TESTING the answer (9 * 10.014084507) / (9 + 10.014084507) = 4.74 90.126760563 / (19.014084507) = 4.74 4.74 = 4.74 Successful !!!